# A Review of LLM-Assisted Ideation

### v2

@@ 1 Introduction
Ideation, the process of generating and developing ideas, is a critical driver of innovation and creativity across diverse domains [osborn1953applied, amabile1988]. Over the years, many techniques and tools have been developed to facilitate ideation, ranging from traditional methods such as brainstorming and the Delphi method to assisted ideation approaches, including machine-driven analytical techniques and idea management systems. However, these approaches often face challenges in generating a wealth of truly novel and valuable ideas, overcoming cognitive biases and fixation effects, and supporting efficient evaluation and selection of potentially many promising ideas [smith2003, blair2007, acar2019]. The emergence of large language models (LLMs) has brought new possibilities for ideation. LLM-assisted ideation has demonstrated its advantages in multiple domains, with LLMs not only being capable of assisting ideation in all stages, but also being widely integrated into systems to fully exploit their potential. Nevertheless, there is currently a lack of research systematically analyzing the stages at which LLMs can assist ideation and how they may do so. Moreover, no studies have yet comprehensively examined how LLMs are integrated into systems to interact with humans. Investigating these aspects can help deepen our understanding of the paradigms of LLM-assisted ideation, facilitate the better use of LLMs to support ideation, and contribute to the development of more efficient ideation systems.
@@ 1.1.1 Traditional Ideation
One of the earliest and most well-known ideation techniques is brainstorming, introduced by Alex Osborn in the 1950s [osborn1953applied]. Brainstorming aims to stimulate creative thinking by encouraging the free generation of a large quantity of ideas in a non-judgmental environment. Since its introduction, brainstorming has been widely adopted and has inspired numerous variations and extensions, such as electronic brainstorming [dennis1993] and brainwriting [heslin2009]. A recent literature review by [maaravi2021ideation] analyzed the advantages and disadvantages of electronic brainstorming and proposed a comprehensive model to support creative ideation.
Other classic ideation techniques include the Delphi method [dalkey1963], which seeks to achieve consensus among a panel of experts through iterative rounds of anonymous idea generation and feedback, and morphological analysis [zwicky1969], which systematically explores the combinatorial space of possible solutions by decomposing a problem into its constituent dimensions. These techniques, along with many others, have been extensively studied and applied in various domains to support ideation [vangundy1988].
However, traditional ideation techniques often face challenges in efficiently navigating large problem spaces, overcoming cognitive biases and fixation effects, generating a wide range of novel ideas, and providing structured support for elaborating and refining ideas [kohn2011collaborative, brem2007innovation].
@@ 1.1.2 Assisted Ideation
In recent years, the rapid advancement of digital technologies has opened up new frontiers for assisted ideation. One notable development is the emergence of machine-driven analytical techniques, such as machine learning, to support idea generation. A comprehensive literature review by [ayele2021systematic] identified and summarized over 20 data-driven and machine-driven techniques for enhancing idea generation, highlighting the potential of these approaches to augment human creativity.
Another significant trend is the development and adoption of idea management systems, which provide digital platforms for capturing, organizing, and managing ideas. A recent literature review by [zhu2023taxonomy] classified and summarized over 30 current idea management tools, underscoring their growing prominence. Idea management systems offer numerous benefits, such as electronically storing ideas to prevent loss and improve accessibility [bakker2006, boeddrich2004, brem2009], enabling ideators to search for and build upon existing ideas [sandstrom2010, stenmark2000, wilson2010, xie2010], and facilitating the submission and management of ideas through standardized templates and categorization schemes [fairbank2003, bothos2008, gamlin2007, vahs2013, xie2010].
Despite these advancements, assisted ideation still faces significant challenges. Machine learning techniques, while powerful, typically require large amounts of training data and may struggle to generate a wealth of truly novel and creative ideas [garfield2001]. While idea management systems have greatly improved the capture and organization of ideas, they often lack advanced capabilities for stimulating idea generation, expanding and refining ideas, and supporting the evaluation and selection of promising ideas [acar2019].
@@ 1.1.3 LLM-assisted Ideation
In recent years, the emergence of large language models (LLMs) has opened up exciting new possibilities for assisted ideation. LLMs, such as GPT-3 [brown2020], are AI systems trained on vast amounts of text data, enabling them to generate human-like text, answer questions, and perform a wide range of language tasks. The potential of LLMs to enhance ideation is immense, as they can serve as powerful tools for generating novel ideas by breaking free from habitual patterns of thought, overcoming fixation effects, and navigating vast problem spaces [smith2003]. Additionally, LLMs can aid in expanding and refining concepts, and even assist in the evaluation and selection of potentially numerous promising ideas by mitigating cognitive biases [maaloe2022, wang2022].
@@ 1.2 Related Work
Study [gerlach2017idea] proposed a conceptual framework for idea management, forming a generic model for enterprise idea management projects. This conceptual framework consists of six main phases: preparation, idea generation, improvement, evaluation, implementation, and deployment. However, their framework only considers the enterprise idea management level, neglecting other groups such as individuals and organizations, and does not consider the conceptual framework at a more general ideation level. Importantly, it was developed before the advent of LLMs and does not consider their unique capabilities and challenges. With the rapidly growing interest in LLMs for ideation, there is currently a lack of frameworks that identify the different stages involved in such processes and guide the development and application of LLM-assisted ideation.
Study [zhu2023taxonomy] classified and summarized current idea management tools and highlighted the importance of studying user interaction and experience with idea management tools. However, their review did not consider the recent emergence of large language models (LLMs) and their potential applications in ideation processes. It is clear that a thorough analysis of how users interact with LLM-assisted ideation systems could provide valuable insights to inform the development of more effective and user-friendly systems. Given the novelty of LLMs and their potential to facilitate ideation, understanding user interaction patterns, preferences, and challenges in the context of LLM-assisted ideation is crucial.
@@ 1.3 Contributions
    We analyze the interaction design and user study of LLM-assisted ideation systems based on the relevant literature. The insights gained from this analysis can serve as valuable guidance for designing and developing more user-friendly and efficient LLM-assisted ideation systems.
@@ 2.3 Screening
To maintain systematic rigor and adherence to established protocols, the evaluation process followed the PRISMA criteria for systematic reviews [moher2009preferred], adapted to incorporate the recommendations proposed by Page et al. [page2021prisma]. The screening process is comprised of the following steps, as depicted in Figure 1. Initially, a comprehensive search across four databases identified 1457 papers, supplemented by an additional 6 papers from Google Scholar, resulting in a total of 1463 records. After removing 439 duplicates, 1024 unique documents were retained for subsequent screening. The first round of screening eliminated 25 books, workshops, or reviews, retaining 999 papers. The second round involved examining titles, abstracts, and keywords to exclude 921 papers unrelated to ”ideation” and ”large language models,” preserving 78 papers. The third round consisted of full-text screening, further excluding 17 papers that did not meet the inclusion criteria, ultimately retaining 61 papers for analysis.
@@ 3.2 Domains and Application Areas
In the design domain, LLMs have been applied to various sub-domains, including general design, product design, game design, visual design, sustainable design, and biological design. These applications encompass a diverse set of downstream tasks, such as project design [kocaballi2023conversational_D1], prototyping [bilgram2023accelerating_D2, ege2024chatgpt_D10], game level design [sudhakaran2023prompt_D3, todd2023level_D4], interior color design [hou2024c2ideas_D13], eco-design [lee2024generating_D7], and protein design [dharuman2023protein_D5].
The writing domain showcases the use of LLMs in general writing tasks [kim2024towards_W2, kim2023cells_W9, reza2024abscribe_W13, benharrak2024writer_W14, yuan2022wordcraft_W5], as well as more specific applications such as prewriting [wan2024felt_W8, petridis2023anglekindling_W6], creative and narrative writing [chung2022talebrush_W3, ghajargar2022redhead_W7, li2024mystery_W11], academic and professional writing [pividori2024publishing_W12, zhang2023visar_W4, goldi2024intelligent_W10], and digital and informal writing [goodman2022lampost_W1, kim2023cells_W9, reza2024abscribe_W13].
[…]
The scientific research domain highlights the application of LLMs in generating research ideas [pu2024ideasynth_S1, liu2024personaflow_S2, radensky2024scideator_S3], as well as in more specific fields such as social psychology [banker2024machine_S4], materials science [chen2024use_S5], and chemical process design [lee2024gpt_S6].
Lastly, the education domain showcases the use of LLMs in instructional design and tools [junior2023chatgpt_D11, huang2023causalmapper_G5, veloso2024forming_D12], academic writing support [goldi2024intelligent_W10], knowledge building [lee2024prompt_G18], and understanding computing concepts [bernstein2024like_G11].
@@
Design General Design project design [kocaballi2023conversational_D1]
conceptual design [wang2023task_D8]
Product design prototyping design [bilgram2023accelerating_D2], [ege2024chatgpt_D10]
interactive devices design [lu2024large_D9]
Game Design game design [lanzi2023chatgpt_D6]
game levels design [sudhakaran2023prompt_D3], [todd2023level_D4]
educational board games design [junior2023chatgpt_D11]
Visual Design interior color design [hou2024c2ideas_D13]
building envelope design [veloso2024forming_D12]
landscape renderings design [huang2024plantography_D14]
Sustainable Design eco-design [lee2024generating_D7]
Biological Design protein design [dharuman2023protein_D5]
Writing General Writing general writing [kim2024towards_W2], [kim2023cells_W9], [reza2024abscribe_W13], [benharrak2024writer_W14], [yuan2022wordcraft_W5]
Prewriting general prewriting [wan2024felt_W8]
develop press release angles [petridis2023anglekindling_W6]
Creative and Narrative Writing story writing [chung2022talebrush_W3], [yuan2022wordcraft_W5], [wan2024felt_W8], [kim2023cells_W9]
fiction writing [ghajargar2022redhead_W7], [benharrak2024writer_W14]
mystery game script writing [li2024mystery_W11]
Academic and Professional Writing scholarly writing [pividori2024publishing_W12], [benharrak2024writer_W14], [yuan2024llmcrit_W15]
argumentative writing [zhang2023visar_W4], [kim2024towards_W2]
student peer review writing [goldi2024intelligent_W10]
report writing [benharrak2024writer_W14]
proposal writing [kim2024towards_W2], [benharrak2024writer_W14]
Digital and Informal Writing email [goodman2022lampost_W1],[kim2023cells_W9],[reza2024abscribe_W13],[benharrak2024writer_W14]
blog post [kim2024towards_W2], [benharrak2024writer_W14], [yuan2024llmcrit_W15]
social media post [reza2024abscribe_W13]
copywriting [kim2023cells_W9]
web article [benharrak2024writer_W14]
newsletter [kim2024towards_W2], [petridis2023anglekindling_W6]
Multimedia Creation 11 Visual Content Generation image generation [cai2023designaid_M1], [wang2023popblends_M2], [brade2023promptify_M3], [chen2024foundation_M8], [kang2024biospark_G14], [huang2024plantography_D14]
image color matching [hou2024c2ideas_D13]
[…]
shape generation [qian2024shape_M5]
icon generation [wu2023iconshop_M7]
Video Editing and Creation video editing [wang2024lave_M4]
Scientific Research General Scientific Research research idea generation [pu2024ideasynth_S1], [liu2024personaflow_S2], [radensky2024scideator_S3]
Social Sciences social psychology hypothesis generation [banker2024machine_S4]
Materials and Experimental Sciences experimental materials hypothesis generation [chen2024use_S5]
Chemical Sciences chemical process design improvement [lee2024gpt_S6]
Education Instructional Design and Tools educational board game design [junior2023chatgpt_D11]
systems thinking education [huang2023causalmapper_G5]
architecture design education [veloso2024forming_D12]
Academic Writing writing process support [goldi2024intelligent_W10]
Knowledge Building knowledge creation process [lee2024prompt_G18]
Knowledge Learning computing concepts understanding [bernstein2024like_G11]
@@ 3.3 General Domain and Ideation Methods
Structured Thinking 1 Thinking Chain Method chain of thought [lee2024prompt_G18], [hou2024c2ideas_D13]
graph of thoughts [besta2024graph_G12]
causal map [huang2023causalmapper_G5]
mind map [liu2024personaflow_S2], [zhang2023visar_W4]
Structured Template template for reflection [xu2024jamplate_G3]
criteria for feedback [yuan2024llmcrit_W15]
principles for ideation [lee2024generating_D7]
conceptual moves for ideation [heyman2024supermind_G2]
Decomposition & Composition Decomposition task-decompose [wang2023task_D8], [qian2024shape_M5]
Composition recombines and mutates [lanzi2023chatgpt_D6]
conceptual blending [wang2023popblends_M2]
combinational creativity [chen2024foundation_M8]
recomposing facts [pu2024ideasynth_S1], [radensky2024scideator_S3]
Analogical Thinking 1 Analogical Thinking analogy mining and generation [bhavya2023cam_G16]
analogy design heuristic [veloso2024forming_D12]
recursion analogies generation [bernstein2024like_G11]
bioinspired analogies and tools [kang2024biospark_G14], [chen2024asknaturenet_G10]
Multi- 11111 Dimensional Thinking 111 Design Space develop design space [suh2024luminate_G13], [cai2023designaid_M1]
Multiple Variations idea variation tool [di2022idea_G4]
multiple writing variations [reza2024abscribe_W13]
Persona-Based Thinking user personas for product design [schuller2024generating_G9]
personas for project design [kocaballi2023conversational_D1]
personas for research ideation [liu2024personaflow_S2]
personas for writing feedback [benharrak2024writer_W14]
Information-driven Thinking Information Recommendation real-time information recommendation [blazevic2024real_G15]
Scenario-Based Thinking 11111 AR-based Scenarios body movement for brainstorming [aikawa2023introducing_G8]
Driving-Based scenarios context-aware for creative incubation [paredes2024creative_G17]
Collaborative Ideation 111 Group Ideation general group ideation [gonzalez2024collaborative_G7], [blazevic2024real_G15], [bilgram2023accelerating_D2], [lanzi2023chatgpt_D6], [ege2024chatgpt_D10], [junior2023chatgpt_D11]
conversational deliberation [rosenberg2023conversational_G6]
brainstorming [aikawa2023introducing_G8]
brainwriting [shaer2024ai_G1]
@@
Structured thinking methods encompass the thinking chain method, which includes generating chains of thought [lee2024prompt_G18, hou2024c2ideas_D13], graphs of thoughts [besta2024graph_G12], causal map [huang2023causalmapper_G5], and mind maps [liu2024personaflow_S2, zhang2023visar_W4]. Additionally, structured templates are utilized to guide ideation, and include those used for reflection [xu2024jamplate_G3], criteria for feedback [yuan2024llmcrit_W15], principles for ideation [lee2024generating_D7], and conceptual moves for ideation [heyman2024supermind_G2].
@@
Decomposition and composition methods involve breaking down tasks into smaller components [wang2023task_D8] and recombining or mutating ideas through various techniques, such as conceptual blending [wang2023popblends_M2], combinational creativity [chen2024foundation_M8], and recomposing facts [pu2024ideasynth_S1, radensky2024scideator_S3].
@@
Analogical thinking methods focus on generating and leveraging analogies for ideation, including analogy mining and generation [bhavya2023cam_G16], analogy design heuristics [veloso2024forming_D12], recursion analogies [bernstein2024like_G11], and bioinspired analogies and tools [kang2024biospark_G14, chen2024asknaturenet_G10].
@@
Multi-dimensional thinking methods encompass developing design spaces [suh2024luminate_G13, cai2023designaid_M1], generating multiple variations of ideas [di2022idea_G4, reza2024abscribe_W13], and employing persona-based thinking for various purposes, such as product design [schuller2024generating_G9], project design [kocaballi2023conversational_D1], research ideation [liu2024personaflow_S2], and writing feedback [benharrak2024writer_W14].
@@
Information-driven thinking involves the use of real-time information recommendation to support ideation [blazevic2024real_G15]. Scenario-based thinking methods include using AR-based scenarios for body movement brainstorming [aikawa2023introducing_G8] and driving-based scenarios for context-aware creative incubation [paredes2024creative_G17].
@@
Lastly, collaborative ideation methods focus on group ideation activities, such as general group ideation [gonzalez2024collaborative_G7, blazevic2024real_G15, bilgram2023accelerating_D2, lanzi2023chatgpt_D6, ege2024chatgpt_D10, junior2023chatgpt_D11], conversational deliberation [rosenberg2023conversational_G6], brainstorming [aikawa2023introducing_G8], and brainwriting [shaer2024ai_G1].

### v3

@@ 1 Introduction
Ideation, the process of generating and developing ideas, is a critical driver of innovation and creativity across diverse domains [1, 2]. Over the years, many techniques and tools have been developed to facilitate ideation, ranging from traditional methods such as brainstorming and the Delphi method to assisted ideation approaches, including machine-driven analytical techniques and idea management systems. However, these approaches often face challenges in generating a wealth of truly novel and valuable ideas, overcoming cognitive biases and fixation effects, and supporting efficient evaluation and selection of potentially many promising ideas [3, 4, 5]. The emergence of large language models (LLMs) has brought new possibilities for ideation. LLM-assisted ideation has demonstrated its advantages in multiple domains, with LLMs not only being capable of assisting ideation in all stages, but also being widely integrated into systems to fully exploit their potential. Nevertheless, there is currently a lack of research systematically analyzing the stages at which LLMs can assist ideation and how they may do so. Moreover, no studies have yet comprehensively examined how LLMs are integrated into systems to interact with humans. Investigating these aspects can help deepen our understanding of the paradigms of LLM-assisted ideation, facilitate the better use of LLMs to support ideation, and contribute to the development of more efficient ideation systems.
@@ 1.1.1 Traditional Ideation
One of the earliest and most well-known ideation techniques is brainstorming, introduced by Alex Osborn in the 1950s [1]. Brainstorming aims to stimulate creative thinking by encouraging the free generation of a large quantity of ideas in a non-judgmental environment. Since its introduction, brainstorming has been widely adopted and has inspired numerous variations and extensions, such as electronic brainstorming [6] and brainwriting [7]. A recent literature review by [8] analysed the advantages and disadvantages of electronic brainstorming and proposed a comprehensive model to support creative ideation.
Other classic ideation techniques include the Delphi method [9], which seeks to achieve consensus among a panel of experts through iterative rounds of anonymous idea generation and feedback, and morphological analysis [10], which systematically explores the combinatorial space of possible solutions by decomposing a problem into its constituent dimensions. These techniques, along with many others, have been extensively studied and applied in various domains to support ideation [11].
However, traditional ideation techniques often face challenges in efficiently navigating large problem spaces, overcoming cognitive biases and fixation effects, generating a wide range of novel ideas, and providing structured support for elaborating and refining ideas [12, 13].
@@ 1.1.2 Assisted Ideation
In recent years, the rapid advancement of digital technologies has opened up new frontiers for assisted ideation. One notable development is the emergence of machine-driven analytical techniques, such as machine learning, to support idea generation. A comprehensive literature review by [14] identified and summarized over 20 data-driven and machine-driven techniques for enhancing idea generation, highlighting the potential of these approaches to augment human creativity.
Another significant trend is the development and adoption of idea management systems, which provide digital platforms for capturing, organizing, and managing ideas. A recent literature review by [15] classified and summarized over 30 current idea management tools, underscoring their growing prominence. Idea management systems offer numerous benefits, such as electronically storing ideas to prevent loss and improve accessibility [16, 17, 18], enabling ideators to search for and build upon existing ideas [19, 20, 21, 22], and facilitating the submission and management of ideas through standardized templates and categorization schemes [23, 24, 25, 26, 22].
Despite these advancements, assisted ideation still faces significant challenges. Machine learning techniques, such as using deep learning to analyse patent documents for generating design concepts [27], are powerful but typically require large training datasets and may struggle to generate truly novel and creative ideas [28]. While idea management systems have greatly improved the capture and organization of ideas, they often lack advanced capabilities for stimulating idea generation, expanding and refining ideas, and supporting the evaluation and selection of promising ideas [5].
@@ 1.1.3 LLM-assisted Ideation
In recent years, the emergence of large language models (LLMs) has opened up exciting new possibilities for assisted ideation. LLMs, such as GPT-3 [29], are AI systems trained on vast amounts of text data, enabling them to generate human-like text, answer questions, and perform a wide range of language tasks. The potential of LLMs to enhance ideation is immense, as they can serve as powerful tools for generating novel ideas by breaking free from habitual patterns of thought, overcoming fixation effects, and navigating vast problem spaces [3]. Additionally, LLMs can aid in expanding and refining concepts, and even assist in the evaluation and selection of potentially numerous promising ideas by mitigating cognitive biases [30, 31].
@@ 1.2 Related Work
Study [32] proposed a conceptual framework for idea management that provides a structured approach to systematically capture, develop, and implement innovative ideas within organizations, forming a generic model for enterprise idea management projects. This conceptual framework consists of six main phases: preparation, idea generation, improvement, evaluation, implementation, and deployment. However, their framework only considers the enterprise idea management level, neglecting other groups such as individuals and organizations, and does not consider the conceptual framework at a more general ideation level. Importantly, it was developed before the advent of LLMs and does not consider their unique capabilities and challenges. With the rapidly growing interest in LLMs for ideation, there is currently a lack of frameworks that identify the different stages involved in such processes and guide the development and application of LLM-assisted ideation.
Study [15] classified and summarized current idea management tools and highlighted the importance of studying user interaction and experience with idea management tools. However, their review did not consider the recent emergence of large language models (LLMs) and their potential applications in ideation processes. It is clear that a thorough analysis of how users interact with LLM-assisted ideation systems could provide valuable insights to inform the development of more effective and user-friendly systems. Given the novelty of LLMs and their potential to facilitate ideation, understanding user interaction patterns, preferences, and challenges in the context of LLM-assisted ideation is crucial.
@@ 1.3 Contributions
    We analyse the interaction design and user study of LLM-assisted ideation systems based on the relevant literature. The insights gained from this analysis can serve as valuable guidance for designing and developing more user-friendly and efficient LLM-assisted ideation systems.
@@ 2.3 Screening
To maintain systematic rigor and adherence to established protocols, the evaluation process followed the PRISMA criteria for systematic reviews [33], adapted to incorporate the recommendations proposed by Page et al. [34]. The screening process is comprised of the following steps, as depicted in Figure 1. Initially, a comprehensive search across four databases identified 1457 papers, supplemented by an additional 6 papers from Google Scholar, resulting in a total of 1463 records. After removing 439 duplicates, 1024 unique documents were retained for subsequent screening. The first round of screening eliminated 25 books, workshops, or reviews, retaining 999 papers. The second round involved examining titles, abstracts, and keywords to exclude 921 papers unrelated to ”ideation” and ”large language models,” preserving 78 papers. The third round consisted of full-text screening, further excluding 17 papers that did not meet the inclusion criteria, ultimately retaining 61 papers for analysis.
@@ 3.2 Domains and Application Areas
In the design domain, LLMs have been applied to various sub-domains, including general design, product design, game design, visual design, sustainable design, and biological design. These applications encompass a diverse set of downstream tasks, such as project design [35], prototyping [36, 37], game level design [38, 39], interior color design [40], eco-design [41], and protein design [42].
The writing domain showcases the use of LLMs in general writing tasks [43, 44, 45, 46, 47], as well as more specific applications such as prewriting [48, 49], creative and narrative writing [50, 51, 52], academic and professional writing [53, 54, 55], and digital and informal writing [56, 44, 45].
In the multimedia creation domain, LLMs have been employed for visual content generation, including image generation [57, 58, 59, 60, 61, 62], picture book generation [63], shape generation [64], and icon generation [65]. Additionally, LLMs have been used for video editing [66].
The scientific research domain highlights the application of LLMs in generating research ideas [67, 68, 69], as well as in more specific fields such as social psychology [70], materials science [71], and chemical process design [72].
Lastly, the education domain showcases the use of LLMs in instructional design and tools [73, 74, 75], academic writing support [55], knowledge building [76], and understanding computing concepts [77].
@@
Design General Design project design [35]
conceptual design [78]
Product design prototyping design [36], [37]
interactive devices design [79]
Game Design game design [80]
game levels design [38], [39]
educational board games design [73]
Visual Design interior color design [40]
building envelope design [75]
landscape renderings design [62]
Sustainable Design eco-design [41]
Biological Design protein design [42]
Writing General Writing general writing [43], [44], [45], [46], [47]
Prewriting general prewriting [48]
develop press release angles [49]
Creative and Narrative Writing story writing [50], [47], [48], [44]
fiction writing [51], [46]
mystery game script writing [52]
Academic and Professional Writing scholarly writing [53], [46], [81]
argumentative writing [54], [43]
student peer review writing [55]
report writing [46]
proposal writing [43], [46]
Digital and Informal Writing email [56],[44],[45],[46]
blog post [43], [46], [81]
social media post [45]
copywriting [44]
web article [46]
newsletter [43], [49]
Multimedia Creation 11 Visual Content Generation image generation [57], [58], [59], [60], [61], [62]
image color matching [40]
picture book generation [63]
shape generation [64]
icon generation [65]
Video Editing and Creation video editing [66]
Scientific Research General Scientific Research research idea generation [67], [68], [69]
Social Sciences social psychology hypothesis generation [70]
Materials and Experimental Sciences experimental materials hypothesis generation [71]
Chemical Sciences chemical process design improvement [72]
Education Instructional Design and Tools educational board game design [73]
systems thinking education [74]
architecture design education [75]
Academic Writing writing process support [55]
Knowledge Building knowledge creation process [76]
Knowledge Learning computing concepts understanding [77]
@@ 3.3 General Domain and Ideation Methods
Structured Thinking 1 Thinking Chain Method chain of thought [76], [40]
graph of thoughts [82]
causal map [74]
mind map [68], [54]
Structured Template template for reflection [83]
criteria for feedback [81]
principles for ideation [41]
conceptual moves for ideation [84]
Decomposition & Composition Decomposition task-decompose [78], [64]
Composition recombines and mutates [80]
conceptual blending [58]
combinational creativity [60]
recomposing facts [67], [69]
Analogical Thinking 1 Analogical Thinking analogy mining and generation [85]
analogy design heuristic [75]
recursion analogies generation [77]
bioinspired analogies and tools [61], [86]
Multi- 11111 Dimensional Thinking 111 Design Space develop design space [87], [57]
Multiple Variations idea variation tool [88]
multiple writing variations [45]
Persona-Based Thinking user personas for product design [89]
personas for project design [35]
personas for research ideation [68]
personas for writing feedback [46]
Information-driven Thinking Information Recommendation real-time information recommendation [90]
Scenario-Based Thinking 11111 AR-based Scenarios body movement for brainstorming [91]
Driving-Based scenarios context-aware for creative incubation [92]
Collaborative Ideation 111 Group Ideation general group ideation [93], [90], [36], [80], [37], [73]
conversational deliberation [94]
brainstorming [91]
brainwriting [95]
@@
Structured thinking methods encompass the thinking chain method, which includes generating chains of thought [76, 40], graphs of thoughts [82], causal map [74], and mind maps [68, 54]. Additionally, structured templates are utilized to guide ideation, and include those used for reflection [83], criteria for feedback [81], principles for ideation [41], and conceptual moves for ideation [84].
@@
Decomposition and composition methods involve breaking down tasks into smaller components [78] and recombining or mutating ideas through various techniques, such as conceptual blending [58], combinational creativity [60], and recomposing facts [67, 69].
@@
Analogical thinking methods focus on generating and leveraging analogies for ideation, including analogy mining and generation [85], analogy design heuristics [75], recursion analogies [77], and bioinspired analogies and tools [61, 86].
@@
Multi-dimensional thinking methods encompass developing design spaces [87, 57], generating multiple variations of ideas [88, 45], and employing persona-based thinking for various purposes, such as product design [89], project design [35], research ideation [68], and writing feedback [46].
@@
Information-driven thinking involves the use of real-time information recommendation to support ideation [90]. Scenario-based thinking methods include using AR-based scenarios for body movement brainstorming [91] and driving-based scenarios for context-aware creative incubation [92].
@@
Lastly, collaborative ideation methods focus on group ideation activities, such as general group ideation [93, 90, 36, 80, 37, 73], conversational deliberation [94], brainstorming [91], and brainwriting [95].